mathslover (mathslover):

Find \(\text{min} \left[ (x_1 - x_2)^2 + (12 + \sqrt{1-{x_1}^2 } - \sqrt{4 x_2} )^2 \right] \) for all \(x_1 , x_2 \in R\)

OpenStudy (anonymous):

@ganeshie8 @mathmate @hartnn @perl

mathslover (mathslover):

I'm myself confused. Not sure whether Calculus will do it or not! But very sure, this has got some relation with the Conic Sections. But again, not sure, how!

OpenStudy (anonymous):

great work :P :D

OpenStudy (anonymous):

I'm thinking \[x_{1} = 1\] so \[\sqrt{1-x_{1}^2} = 0\] and you get something that is easier to plot.

OpenStudy (perl):

you want to minimize f(x,y)

OpenStudy (anonymous):

On that plot, you'll see a min value

OpenStudy (perl):

We calculate the Hessian

OpenStudy (perl):

it should come out the same as what ganeshie8 put into wolfram

OpenStudy (perl):

min {(x1-x2)^2+(12+sqrt(1-x1^2)-sqrt(4*x2))^2} = 1/3 (433-7775/(140112 sqrt(2919)-7538831)^(1/3)+(140112 sqrt(2919)-7538831)^(1/3))

OpenStudy (perl):

which is approximately 72.41145427

OpenStudy (anonymous):

\((x_1-x_2)^2\geq 0\) \((12+\sqrt{1-x_1^2}-\sqrt{4x_2})^2\geq0\) ----------------------------------------- the sum \(\geq\) 0 so, the minimum if there is some is =0

mathslover (mathslover):

I don't think that there are any possible values for the function = 0 .

OpenStudy (anonymous):

\(x_1-x_2\geq \pm (12+\sqrt{1-x_1^2}-\sqrt{4x_2})\)

OpenStudy (perl):

First find and classify the critical points of f(x,y) = [ (x - y)^2 + (12 + sqrt(1-x^2) - sqrt(4y) )^2 ] Solve the system of equations fx = 0 fy = 0

mathslover (mathslover):

Finding the critical points... hmm, but how?

OpenStudy (perl):

fx = partial f(x,y) / dx fy = partial f(x,y) /dy

OpenStudy (perl):

Paul explains it here http://tutorial.math.lamar.edu/Classes/CalcIII/RelativeExtrema.aspx

mathslover (mathslover):

Okay so here is what I can think about : Let : \( 12 + \sqrt{1 - {x_1}^2 } = A \) \((A-12)^2 = 1 - {x_1} ^2\) And say, \(\sqrt{4 x_2} = B\) \(B^2= 4 x_2 \) Which is infact an equation of parabola.

mathslover (mathslover):

Trust me, it is an objective question and as per the time given for each question (2-3 minutes), I don't think that doing this question with calculus will be appropriate. It has really got something to do with Conics. @Kainui @vishweshshrimali5

OpenStudy (anonymous):

Are you allowed to use some graphing tool?

mathslover (mathslover):

Unfortunately, no @aleroth . Only allowed to use ... one pen, paper.. .and that's it :(

OpenStudy (perl):

ok, so you are looking for a pre-calculus argument

Parth (parthkohli):

Is this question listed only under "conic sections" or some specific chapter under it? I haven't really started with conics but one geometric interpretation, which I'm not sure is a good start, would be considering these two points:\[P(x_1, 12 + \sqrt{1 - x^2})\]\[Q(x_2, 2\sqrt{x_2})\]...and then minimizing the distance between them.

Parth (parthkohli):

Now if we do find some nice way to minimize distances other than using the distance formula, then this problem would be a piece of cake. But the reason why this problem is a good one is that I don't think there is. :P

mathslover (mathslover):

Exactly! That is what I was trying to do. Haha! kya karun yaar mujhe buddha mil gaya... :/ This question is seriously... :'( Mind boggling :P

ganeshie8 (ganeshie8):

Very interesting! so you're basically trying to find the minimum distance between a circle and a parabola : \(x^2+(y-12)^2 = 1\) \(y^2=4x\)

ganeshie8 (ganeshie8):

those equations follow trivially from PK's previous reply ^

Parth (parthkohli):

Oh wow, I somehow skipped mathslover's reply. That looks like it.

ganeshie8 (ganeshie8):

Is there an easy way to find the distance between conics ? XD

Parth (parthkohli):

I cheated a little... https://answers.yahoo.com/question/index?qid=20120203010806AAxz86W

mathslover (mathslover):

I'm going to exile... for 100 years.. LOL!

OpenStudy (perl):

it would be nice to see if you get the same answer as the calculus solution :)

OpenStudy (perl):

and how did wolfram get the answer so quickly (did they use calculus)?

OpenStudy (kainui):

A trick I learned in calculus 1 for minimzation problems is that you can look for things that don't affect the minimum point and remove them, such as square roots or constants. Not sure if that'll help here or not, but probably. A good example would be: \[\Large f(x) = \sqrt{1-x^2} \\ \Large g(x) = 1-x^2\] both have the same minimum for x. Another example: \[\Large f(x)= x^2+7 \\ \Large g(x) = x^2\] Also have the same minimum for x. So some things to think about to quickly throw away some stuff maybe?

OpenStudy (kainui):

*Just be careful when using this rule ;)

ganeshie8 (ganeshie8):

this question made my day! working it is a simple matter of finding the normal and yes it can be done in less than 2 minutes XD

OpenStudy (kainui):

Another thing (didn't read so might have been said) x2 has to be greater than 0 otherwise you get a negative under the square root.

OpenStudy (perl):

Using this widget I get the critical point (x,y) is (-sqrt(5)/5, 4) so the minimum is f( -sqrt(5)/5, 4) http://www.wolframalpha.com/widgets/view.jsp?id=597dd91e6b177f393877f173f231593f

OpenStudy (kainui):

Woah this geometric interpretation seems like the way to go, but I don't think I would have gotten there on my own haha.

OpenStudy (perl):

yes

ganeshie8 (ganeshie8):

this is reallly a very seductive trick! but it only works in this special case and im sure im gona forget this after few days

ganeshie8 (ganeshie8):

the author cooked up the problem by taking two conics and squaring the distance between them - what cunning mind haha!

OpenStudy (kainui):

Can someone help me draw this picture, I'm kind of stuck: